# The Leftmost particle of branching subordinators

### v1

@@ 3.2.2 The case $r<\rho$
We assume $r<\rho$. Let us fix $k\geq 0$ and note from the proof of Lemma 3.3 that $\mathbb{P}(\underline{Y}(\infty)=0)>0$ so $\mathbb{P}(\underline{Y}(\infty)\geq k+1)<1$. Decomposing on the first transition of the system (we recall that the initial configuration is a single particle located at the position 00) and using the branching property we get
@@
If the binary branchings from the branching Poisson process are replaced by branchings of order $n$ for some $n\geq 3$, then obtaining a similar recursive relation for the law of $\underline{Y}(\infty)$ requires to solve a polynomial equation of order $n$. This suggests that, in general, there is no simple expression for the law of $\underline{Y}(\infty)$, even in the case of branching subordinators supported on $\mathbb{Z}_{+}$.
@@ 4.1 Lower bound
Let $C$ be as in Assumption 1 and $c_{1}$ be as in Assumption 4. Set $\gamma$ as in Theorem 1.5 and $m_{0}:=(C/c_{1}^{1-\alpha})^{\gamma}$. For any $m\in(0,m_{0})$ we have
@@
  • Proof.
    We recall that, for any $a>a_{0}(\Lambda)$ (see Section 1.2), $Y^{a}$ denotes the $a$-truncation of $Y$, described in Section 1.2 and rigorously introduced in Definition A.1. Let us fix $m\in(0,m_{0})$ and $t>(a_{0}(\Lambda)/m)^{1/\gamma}$. Reasoning similarly as in the proof of Lemma 3.2, we get $\{\underline{Y}(t)<mt^{\gamma}\}=\{\underline{Y}^{mt^{\gamma}}(t)<mt^{\gamma}\}$ so, for any $\lambda\geq 0$, we have
@@
    which is nothing but $e^{\lambda mt^{\gamma}-t\kappa_{mt^{\gamma}}(\lambda)}$ by (46) (see also Section 1.2 where (46) is announced). Combining the above with (10) we get
@@ 4.2 Upper bound
Fix $m>0$ and $q>0$ that will be determined later and $k_{0}\geq 1$ such that $m2^{k_{0}\gamma-1}>a_{0}(\Lambda)$ (see Section 1.2). We choose a $k\geq k_{0}$ and set $a:=m2^{k\gamma-1}$. We consider $Y^{a}$. We denote by $\mathscr{U}_{k}$ the set of indices of particles that are born on the time interval $[2^{k-1},2^{k})$, ie $\mathscr{U}_{k}:=\{u\in\mathscr{U},\mathfrak{b}^{a}_{u}\in[2^{k-1},2^{k})\}$ (see Appendix A for the definition of the set of indices $\mathscr{U}$ and the birth time $\mathfrak{b}^{a}_{u}$). We denote by $S_{br}^{a}$ the set of times at which particles of $Y^{a}$ experience branchings in $\mathscr{Q}^{\star}$ (see (41) or after (3)). For $t\in S_{br}^{a}$ we denote by $u(t)\in\mathscr{U}$ the index of the particle branching at time $t$ and by $\bm{x}(t)=(x_{i}(t))_{i\geq 1}$ the corresponding element of $\mathscr{Q}^{\star}$, that is, $x_{1}(t)=Y^{a}_{u(t)}(t)-Y^{a}_{u(t)}(t-)$ and for any $i\geq 1$, $x_{i+1}(t)$ is the distance between $Y^{a}_{u(t)}(t-)$ and the birth position of the $i$-th child of $u(t)$ at time $t$. We define the random sets
$S_{low,k}$ :=$:=\{t\in[2^{k},2^{k+1})\cap S^{a}_{br},u(t)\in\mathscr{U}_{k},Y^{a}_{u(t)}(t-)=\min_{v\in\mathscr{U}_{k}}Y^{a}_{v}(t-)\},$
$\mathscr{D}_{k}$ :=$:=\{(t,i,\mathfrak{z}):t\in S_{low,k},i\geq 2,\mathfrak{z}=x_{i}(t)<\infty\}.$
The set $\mathscr{D}_{k}$ contains information on branchings during the time interval $[2^{k},2^{k+1})$, where the branching particle is the leftmost one among those that are born on the time interval $[2^{k-1},2^{k})$. Lemma A.7 shows that the minimum in the definition of $\mathscr{D}_{k}$ is almost surely reached by exactly one particle. Using Lemma A.7, the branching property and that all trajectories of particles are càd-làg, it is not difficult to see that, conditionally on $\{\sharp\mathscr{U}_{k}\geq 1\}$, $\min S_{low,k}\sim 2^{k}+Z\mathds{1}_{Z<2^{k}}+\infty\mathds{1}_{Z\geq 2^{k}}$ where $Z$ is exponentially distributed with parameter $\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a\})<\infty$ (recall from (4) that $\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a\})<\infty$). By the strong branching property (see Remark A.6), we deduce that $\{(t,\bm{x}(t)),\ t\in S_{low,k}\}$ is a Poisson point process on $[2^{k},2^{k+1})\times\mathscr{Q}^{\star}$ with intensity measure $\mathrm{d}t\otimes\Lambda(\cdot\cap\{\bm{x}\in\mathscr{Q},x_{2}<a\})$. In particular, $S_{low,k}$ is a discrete subset of $[2^{k},2^{k+1})$. We have from the above and (4) that $\mathbb{E}[\sharp\mathscr{D}_{k}]<\infty$ so $\mathscr{D}_{k}$ is almost surely finite. We then define
@@
$B_{k}$ :=$:=\{S_{k}\geq 2^{k-1}M_{m2^{k\gamma-1}}(q2^{k\gamma\sigma})\}.$ (29)
For any $(t,i,\mathfrak{z})\in\mathscr{D}_{k}$, introduce $\bm{\xi}_{t,i,\mathfrak{z}}:=\bm{\xi}^{a}_{u_{i}(t)}$, where $u_{i}(t)$ is the $(i-1)^{th}$ child born from the particle $u(t)$ at time $t$ (the child born with displacement $x_{i}(t)=\mathfrak{z}$) and where we recall that, for a particle $v$, $\bm{\xi}^{a}_{v}$ is the canonical trajectory (associated with $Y^{a}$) of the particle $v$, see below (3) for a description of canonical trajectories and Definition A.2 for a rigorous definition. In particular, we see from the construction in Appendix A that, conditionally on $\mathscr{D}_{k}$, $(\bm{\xi}_{t,i,z})_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}$ is a collection of i.i.d copies of $(\bm{\xi}(t);t\geq 0)$, a subordinator with Laplace exponent $\phi$ and starting position 00. For any $k\geq k_{0}$ we let
$C_{k}:=\{\min\{\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2}),(t,i,\mathfrak{z})\in\mathscr{D}_{k}\}\leq m2^{k\gamma}\},$ (30)
@@
  • Proof.
    Fix $k\geq k_{0}$ and set $a:=m2^{k\gamma-1}$ as before. Define $F_{k}:\mathscr{Q}^{\star}\rightarrow\mathbb{R}_{+}$ by $F_{k}(\bm{x}):=\sum_{i\geq 2}e^{-x_{i}q2^{k\gamma\sigma}}\mathds{1}_{\{x_{i}<a\}}$ and note that $S_{k}=\sum_{t\in S_{low,k}}F_{k}(\bm{x}(t))$. Recall that, conditionally on $\{\sharp\mathscr{U}_{k}\geq 1\}$, $\{(t,\bm{x}(t)),\ t\in S_{low,k}\}$ is a Poisson point process on $[2^{k},2^{k+1})\times\mathscr{Q}^{\star}$ with intensity measure $\mathrm{d}t\otimes\Lambda(\cdot\cap\{\bm{x}\in\mathscr{Q},x_{2}<a\})$. In particular, if we let $(X_{k}^{i})_{i\geq 1}$ be i.i.d random variables distributed as $F_{k}(\bm{x})$ under the law $\Lambda(\cdot\cap\{\bm{x}\in\mathscr{Q},x_{2}<a\})/\Lambda(\{\bm{x}\in\mathscr{Q},x_{2}<a\})$ on $\mathscr{Q}^{\star}$, and $N_{k}$ be a Poisson random variable with parameter $2^{k}\Lambda(\{\bm{x}\in\mathscr{Q},x_{2}<a\})$, with $\{\sharp\mathscr{U}_{k}\geq 1\}$, $N_{k}$ and $(X_{k}^{i})_{i\geq 1}$ being independent, then $S_{k}$ is distributed as $\mathds{1}_{\sharp\mathscr{U}_{k}\geq 1}\tilde{S}_{k}$ with $\tilde{S}_{k}:=\sum_{j=1}^{N_{k}}X_{k}^{j}$. Using classical identities for compound Poisson distributions and the definitions of $M_{\cdot}(\cdot)$ and $M_{2,\cdot}(\cdot)$ in (11) and (18) we get
@@
    $\mathbb{P}(B_{k}^{c})$ $\leq\mathbb{P}(\tilde{S}_{k}\leq\mathbb{E}[\tilde{S}_{k}]/2)+\mathbb{P}(\sharp\mathscr{U}_{k}=0)\leq\mathbb{P}(|S_{k}-\mathbb{E}[S_{k}]|\geq\mathbb{E}[S_{k}]/2)+\mathbb{P}(\sharp\mathscr{U}_{k}=0)$
    $\leq\frac{4Var(S_{k})}{\mathbb{E}[S_{k}]^{2}}+\mathbb{P}(\sharp\mathscr{U}_{k}=0)=2^{-k}\frac{4M_{2,m2^{k\gamma-1}}(q2^{k\gamma\sigma})}{(M_{m2^{k\gamma-1}}(q2^{k\gamma\sigma}))^{2}}+\mathbb{P}(\sharp\mathscr{U}_{k}=0).$
    Since the initial particle $\varnothing$ of $Y^{a}$ gives birth to new particles at rate $\Lambda(\{\bm{x}\in\mathscr{Q},x_{2}<a\})\geq\Lambda(\{\bm{x}\in\mathscr{Q},x_{2}<m2^{k_{0}\gamma-1}\})>0$ we get $\mathbb{P}(\sharp\mathscr{U}_{k}=0)\leq e^{-2^{k-1}\Lambda(\{\bm{x}\in\mathscr{Q},x_{2}<m2^{k_{0}\gamma-1}\})}$. Since $2^{\sigma}q/m^{\sigma}\leq c_{0}$, (19) from Assumption 3 holds true with $c=2^{\sigma}q/m^{\sigma}$. We get that the above is, for large $k$, bounded by $2^{-k}L$ for some constant $L>0$. This yields $\sum_{k\geq k_{0}}\mathbb{P}(B_{k}^{c})<\infty$. ∎
@@
  • Proof.
    Recall that $\bm{\xi}$ denotes a subordinator with Laplace exponent $\phi$ and starting position 00. We let ${P}(\cdot)$ be the probability measure associated with $\bm{\xi}$ and ${E}[\cdot]$ the associated expectation. For $k\geq k_{0}$ we have
@@
    We now study $\sum_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}{P}(\mathfrak{z}+\bm{\xi}(2^{k+2})\leq m2^{k\gamma})$ using a change of measure argument that is used in the proof of Theorem III.11 of [4] and that comes from large deviations theory. For $\omega\in(0,\infty)$, let ${P}^{(\omega)}(\cdot)$ be defined by ${P}^{(\omega)}(A)={E}\left[e^{-\omega\bm{\xi}(t)+\phi(\omega)t}\mathds{1}_{A}\right]$ for $A\in\mathscr{F}_{t}$ and $t\geq 0$, where $\mathscr{F}_{t}$ is the sigma field generated by $(\bm{\xi}(s))_{s\in[0,t]}$. We see from Section III.4 of [4] that this definition is consistent and that, under ${P}^{(\omega)}(\cdot)$, $\bm{\xi}$ is a subordinator with Laplace exponent $\phi^{(\omega)}(\cdot):=\phi(\omega+\cdot)-\phi(\omega)$ so, in particular,
@@
    ×$\times{P}^{(\omega)}\left((m2^{k\gamma}-\mathfrak{z})/2\leq\bm{\xi}(2^{k+2})\leq m2^{k\gamma}-\mathfrak{z}\right).$ (33)
@@
    where we have used Chebyshev inequality and (32). From Assumption 1 and the monotone density theorem we have that $\phi^{\prime}(\lambda)\sim\alpha C\lambda^{\alpha-1}$ and $\phi^{\prime\prime}(\lambda)\sim-\alpha(1-\alpha)C\lambda^{\alpha-2}$ as $\lambda$ goes to infinity so, in particular, $(\phi^{\prime})^{-1}(r)\sim(\alpha C/r)^{1/(1-\alpha)}$ as $r$ goes to zero. We get that $-\phi^{\prime\prime}((\phi^{\prime})^{-1}(r))\sim C^{\prime}r^{(2-\alpha)/(1-\alpha)}$ (where $C^{\prime}:=(1-\alpha)(\alpha C)^{-1/(1-\alpha)}$) so there is $r_{0}>0$ such that $-\phi^{\prime\prime}((\phi^{\prime})^{-1}(r))\leq 2C^{\prime}r^{(2-\alpha)/(1-\alpha)}$ for all $r\leq r_{0}$. We choose $k_{1}(m)\geq k_{0}(m)$ such that $m2^{k\gamma}/2^{k+2}\leq r_{0}$ for all $k\geq k_{1}(m)$. We get, whenever $k\geq k_{1}(m)$, for any $\mathfrak{z}\in(0,m2^{k\gamma-1})$, that $-\phi^{\prime\prime}(\omega_{\mathfrak{z}})\leq 2C^{\prime}(3\times 2^{-(k+4)}(m2^{k\gamma}-\mathfrak{z}))^{(2-\alpha)/(1-\alpha)}$. In this case, the last term in (34) is larger than
@@
  • Proof of (27).
    Since $C_{k}\subset\{\mathscr{D}_{k}\neq\emptyset\}$, we have almost surely that, on $C_{k}$, $\{\underline{Y}^{m2^{k\gamma-1}}(t-)+\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2}-t),(t,i,\mathfrak{z})\in\mathscr{D}_{k}\}$ is a non-empty subset of the set of positions of particles in the system $Y^{m2^{k\gamma-1}}$ at time $2^{k+2}$. In particular, we have almost surely that, on $C_{k}$,
    $\underline{Y}^{m2^{k\gamma-1}}(2^{k+2})$ $\leq\min_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}\{\underline{Y}^{m2^{k\gamma-1}}(t-)+\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2}-t)\}\leq\underline{Y}^{m2^{k\gamma-1}}(2^{k+1})+\min_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}\{\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2})\}$
    $\leq\underline{Y}^{m2^{k\gamma-1}}(2^{k+1})+m2^{k\gamma}\leq\underline{Y}^{m2^{(k-1)\gamma-1}}(2^{k+1})+m2^{k\gamma}.$ (40)
@@
    $\underline{Y}(s)\leq\underline{Y}(2^{k+1})\leq\underline{Y}^{m2^{(k-1)\gamma-1}}(2^{k+1})\leq\underline{Y}^{m2^{(K_{0}-1)\gamma-1}}(2^{K_{0}+1})+\sum_{j=K_{0}}^{k-1}m2^{j\gamma}.$
@@
    $\sup_{s\in[2^{k},2^{k+1}]}\frac{\underline{Y}(s)}{s^{\gamma}}\leq\frac{m}{2^{\gamma}-1}+2^{-k\gamma}\left(\underline{Y}^{m2^{(K_{0}-1)\gamma-1}}(2^{K_{0}+1})-\frac{m2^{K_{0}\gamma}}{2^{\gamma}-1}\right).$

### v2

@@ 3.2.2 The case $r<\rho$
We assume $r<\rho$. Let us fix $k\geq 0$ and note from the proof of Lemma 3.3 that $\mathbb{P}(\underline{Y}(\infty)=0)>0$ so $\mathbb{P}(\underline{Y}(\infty)\geq k+1)<1$. Decomposing on the first transition of the system (we recall that the initial configuration is a single particle located at the position 0) and using the branching property we get
@@
If the binary branchings from the branching Poisson process are replaced by $p$-ary branchings for some $p\geq 3$, then obtaining a similar recursive relation for the law of $\underline{Y}(\infty)$ requires to solve a polynomial equation of degree $p$. This suggests that, in general, there is no simple expression for the law of $\underline{Y}(\infty)$, even in the case of branching subordinators supported on $\mathbb{Z}_{+}$.
@@ 4.1 Lower bound
Let $\alpha,C$ be as in Assumption 1 and $\sigma,c_{1}$ be as in Assumption 5. Set $\gamma$ as in Theorem 1.5 and $m_{0}:=(C/c_{1}^{1-\alpha})^{\gamma}$. For any $m\in(0,m_{0})$ we have
@@
We recall that, for any $a>a_{0}(\Lambda)$ (see Section 1.2), $Y^{a}$ denotes the $a$-truncation of $Y$, described in Section 1.2 and rigorously introduced in Definition A.1. Let us fix $m\in(0,m_{0})$ and $t>(a_{0}(\Lambda)/m)^{1/\gamma}$. Reasoning similarly as in the proof of Lemma 3.2, we get $\{\underline{Y}(t)<mt^{\gamma}\}=\{\underline{Y}^{mt^{\gamma}}(t)<mt^{\gamma}\}$ (where $Y^{mt^{\gamma}}$ is $Y^{a}$ with the choice $a=mt^{\gamma}$) so, for any $\lambda\geq 0$, we have
@@
which is nothing but $e^{\lambda mt^{\gamma}-t\kappa_{mt^{\gamma}}(\lambda)}$ by (13). Combining the above with (11) we get
@@ 4.2 Upper bound
Fix $m>0$ and $q>0$ that will be determined later and $k_{0}\geq 1$ such that $m2^{k_{0}\gamma-1}>a_{0}(\Lambda)$ (see Section 1.2). We choose a $k\geq k_{0}$ and set $a=a(k):=m2^{k\gamma-1}$. We consider $Y^{a}$ (as described in Section 1.2 and rigorously introduced in Definition A.1) with this choice of $a$. We denote by $\mathscr{U}_{k}^{a}$ the set of indices of particles that are born on the time interval $2^{k-1},2^{k})$, ie $\mathscr{U}_{k}^{a}:=\{u\in\mathscr{U};\mathfrak{b}^{a}_{u}\in[2^{k-1},2^{k})\}$ (see Appendix A for the definition of the set of indices $\mathscr{U}$ and the birth time $\mathfrak{b}^{a}_{u}$). We denote by $S_{br}^{a}$ the set of times at which particles of $Y^{a}$ experience branchings in $\mathscr{Q}^{\star}$ (see (44) or after (3)). For $t\in S_{br}^{a}$ we denote by $u(t)\in\mathscr{U}$ the index of the particle branching at time $t$ and by $\bm{x}(t)=(x_{i}(t))_{i\geq 1}$ the corresponding element of $\mathscr{Q}^{\star}$, that is, $x_{1}(t)=Y^{a}_{u(t)}(t)-Y^{a}_{u(t)}(t-)$ and for any $i\geq 1$, $x_{i+1}(t)$ is the distance between $Y^{a}_{u(t)}(t-)$ and the birth position of the $i$-th child of $u(t)$ at time $t$. We define the random sets
$S_{low,k}$ $:=\{t\in[2^{k},2^{k+1})\cap S^{a}_{br};u(t)\in\mathscr{U}_{k}^{a},Y^{a}_{u(t)}(t-)=\min_{v\in\mathscr{U}_{k}^{a}}Y^{a}_{v}(t-)\},$
$\mathscr{D}_{k}$ $:=\{(t,i,\mathfrak{z});t\in S_{low,k},i\geq 2,\mathfrak{z}=x_{i}(t)<\infty\}.$
The set $\mathscr{D}_{k}$ contains information on branchings during the time interval $2^{k},2^{k+1})$, where the branching particle is the leftmost one among those that are born on the time interval $2^{k-1},2^{k})$. Lemma A.7 shows that the minimum in the definition of $\mathscr{D}_{k}$ is almost surely reached by exactly one particle. The following lemma relies on Assumption 4 and tells us that the branchings recorded in $S_{low,k}$ and $\mathscr{D}_{k}$ occur similarly to branchings of a generic particle. Recall the measure $\Lambda^{a}(\mathrm{d}\bm{x})$ from Section 1.2.
Lemma 4.2 (branchings of the leftmost particle).
Conditionally on $\{\sharp\mathscr{U}_{k}^{a}\geq 1\}$, the point process $\{(t,\bm{x}(t));\ t\in S_{low,k}\}$ on $2^{k},2^{k+1})\times\mathscr{Q}^{\star}$ is a Poisson point process with intensity measure $\mathrm{d}t\otimes\Lambda^{a}(\cdot\cap\mathscr{Q}^{\star})$.
The proof of Lemma 4.2 is straightforward and shifted to the end of Appendix A. Since $\Lambda^{a}(\mathscr{Q}^{\star})=\Lambda(\{\bm{x}\in\mathscr{Q};\;x_{2}<a\})<\infty$ by (4) and, by (4) again, the average number of new particles arriving in $Y^{a}$ at each jumping time of the above point process is finite. Therefore, $\mathbb{E}[\sharp\mathscr{D}_{k}]<\infty$ so $\mathscr{D}_{k}$ is almost surely finite (in particular, $S_{low,k}$ is a discrete subset of $2^{k},2^{k+1})$). We then define
@@
$B_{k}$ $:=\{S_{k}\geq 2^{k-1}M_{m2^{k\gamma-1}}(q2^{k\gamma\sigma})\}.$ (32)
For any $(t,i,\mathfrak{z})\in\mathscr{D}_{k}$, introduce $\bm{\xi}_{t,i,\mathfrak{z}}:=\bm{\xi}^{a}_{u_{i}(t)}$, where $u_{i}(t)$ is the $(i-1)^{th}$ child born from the particle $u(t)$ at time $t$ (the child born with displacement $x_{i}(t)=\mathfrak{z}$) and where we recall that, for a particle $v$, $\bm{\xi}^{a}_{v}$ is the canonical trajectory (associated with $Y^{a}$) of the particle $v$, see below (3) for a description of canonical trajectories and Definition A.2 for a rigorous definition. In particular, we see from the construction in Appendix A that, conditionally on $\mathscr{D}_{k}$, $(\bm{\xi}_{t,i,z})_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}$ is a collection of i.i.d copies of $(\bm{\xi}(t);t\geq 0)$, a subordinator with Laplace exponent $\phi$ and starting position 0. For any $k\geq k_{0}$ we let
$C_{k}:=\{\min\{\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2});(t,i,\mathfrak{z})\in\mathscr{D}_{k}\}\leq m2^{k\gamma}\},$ (33)
@@
Fix $k\geq k_{0}$ and set $a:=m2^{k\gamma-1}$ as before. Define $F_{k}:\mathscr{Q}^{\star}\rightarrow\mathbb{R}_{+}$ by $F_{k}(\bm{x}):=\sum_{i\geq 2}e^{-x_{i}q2^{k\gamma\sigma}}\mathds{1}_{\{x_{i}<a\}}$ and note that $S_{k}=\sum_{t\in S_{low,k}}F_{k}(\bm{x}(t))$. We know from Lemma 4.2 that, conditionally on $\{\sharp\mathscr{U}_{k}^{a}\geq 1\}$, $\{(t,\bm{x}(t));\ t\in S_{low,k}\}$ is a Poisson point process on $2^{k},2^{k+1})\times\mathscr{Q}^{\star}$ with intensity measure $\mathrm{d}t\otimes\Lambda^{a}(\cdot\cap\mathscr{Q}^{\star})$. In particular, if we let $(X_{k}^{i})_{i\geq 1}$ be i.i.d random variables distributed as $F_{k}(\bm{z})$ (where $\bm{z}$ has distribution $\Lambda^{a}(\cdot\cap\mathscr{Q}^{\star})/\Lambda^{a}(\mathscr{Q}^{\star})$ on $\mathscr{Q}^{\star}$), and $N_{k}$ be a Poisson random variable with parameter $2^{k}\Lambda^{a}(\mathscr{Q}^{\star})$, with $\{\sharp\mathscr{U}_{k}^{a}\geq 1\}$, $N_{k}$ and $(X_{k}^{i})_{i\geq 1}$ being independent, then $S_{k}$ is distributed as $\mathds{1}_{\sharp\mathscr{U}_{k}^{a}\geq 1}\tilde{S}_{k}$ with $\tilde{S}_{k}:=\sum_{j=1}^{N_{k}}X_{k}^{j}$. Using classical identities for compound Poisson distributions and the definitions of $M_{\cdot}(\cdot)$ and $M_{2,\cdot}(\cdot)$ in (12) and (22) we get
@@
$\mathbb{P}(B_{k}^{c})$ $\leq\mathbb{P}(\tilde{S}_{k}\leq\mathbb{E}[\tilde{S}_{k}]/2)+\mathbb{P}(\sharp\mathscr{U}_{k}^{a}=0)\leq\mathbb{P}(|S_{k}-\mathbb{E}[S_{k}]|\geq\mathbb{E}[S_{k}]/2)+\mathbb{P}(\sharp\mathscr{U}_{k}^{a}=0)$
$\leq\frac{4Var(S_{k})}{\mathbb{E}[S_{k}]^{2}}+\mathbb{P}(\sharp\mathscr{U}_{k}^{a}=0)=2^{-k}\frac{4M_{2,m2^{k\gamma-1}}(q2^{k\gamma\sigma})}{(M_{m2^{k\gamma-1}}(q2^{k\gamma\sigma}))^{2}}+\mathbb{P}(\sharp\mathscr{U}_{k}^{a}=0).$
Since the initial particle $\varnothing$ of $Y^{a}$ gives birth to new particles at rate $\Lambda^{a}(\mathscr{Q}^{\star})=\Lambda(\{\bm{x}\in\mathscr{Q};x_{2}<a\})\geq\Lambda(\{\bm{x}\in\mathscr{Q};x_{2}<m2^{k_{0}\gamma-1}\})>0$ we get $\mathbb{P}(\sharp\mathscr{U}_{k}^{a}=0)\leq e^{-2^{k-1}\Lambda(\{\bm{x}\in\mathscr{Q};x_{2}<m2^{k_{0}\gamma-1}\})}$. Since $2^{\sigma}q/m^{\sigma}\leq c_{0}$, (23) from Assumption 3 holds true with $c=2^{\sigma}q/m^{\sigma}$. We get that the above is, for large $k$, bounded by $2^{-k}L$ for some constant $L>0$. This yields $\sum_{k\geq k_{0}}\mathbb{P}(B_{k}^{c})<\infty$. ∎
@@
Recall that $\bm{\xi}$ denotes a subordinator with Laplace exponent $\phi$ and starting position 0. We let ${P}(\cdot)$ be the probability measure associated with $\bm{\xi}$ and ${E}[\cdot]$ the associated expectation. For $k\geq k_{0}$ we have
@@
We now study $\sum_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}{P}(\mathfrak{z}+\bm{\xi}(2^{k+2})\leq m2^{k\gamma})$ using a change of measure argument that is used in the proof of Theorem III.11 of [5] and that comes from large deviations theory. For $\omega\in(0,\infty)$, let ${P}^{(\omega)}(\cdot)$ be defined by ${P}^{(\omega)}(A):={E}\left[e^{-\omega\bm{\xi}(t)+\phi(\omega)t}\mathds{1}_{A}\right]$ for $A\in\mathscr{F}_{t}$ and $t\geq 0$, where $\mathscr{F}_{t}$ is the sigma field generated by $(\bm{\xi}(s))_{s\in[0,t]}$. We see from Section III.4 of [5] that this definition is consistent and that, under ${P}^{(\omega)}(\cdot)$, $\bm{\xi}$ is a subordinator with Laplace exponent $\phi^{(\omega)}(\cdot):=\phi(\omega+\cdot)-\phi(\omega)$ so, in particular,
@@
$\times{P}^{(\omega)}\left((m2^{k\gamma}-\mathfrak{z})/2\leq\bm{\xi}(2^{k+2})\leq m2^{k\gamma}-\mathfrak{z}\right).$ (36)
@@
where we have used Chebyshev inequality and (35). From Assumption 1 and the monotone density theorem we have that $\phi^{\prime}(\lambda)\sim\alpha C\lambda^{\alpha-1}$ and $\phi^{\prime\prime}(\lambda)\sim-\alpha(1-\alpha)C\lambda^{\alpha-2}$ as $\lambda$ goes to infinity so, in particular, $(\phi^{\prime})^{-1}(r)\sim(\alpha C/r)^{1/(1-\alpha)}$ as $r$ goes to zero. We get that $\phi^{\prime\prime}((\phi^{\prime})^{-1}(r))\sim C^{\prime}r^{(2-\alpha)/(1-\alpha)}$ (where $C^{\prime}:=(1-\alpha)(\alpha C)^{-1/(1-\alpha)}$) so there is $r_{0}>0$ such that $\phi^{\prime\prime}((\phi^{\prime})^{-1}(r))\leq 2C^{\prime}r^{(2-\alpha)/(1-\alpha)}$ for all $r\leq r_{0}$. We choose $k_{1}(m)\geq k_{0}(m)$ such that $m2^{k\gamma}/2^{k+2}\leq r_{0}$ for all $k\geq k_{1}(m)$. We get, whenever $k\geq k_{1}(m)$, for any $\mathfrak{z}\in(0,m2^{k\gamma-1})$, that $\phi^{\prime\prime}(\omega_{\mathfrak{z}})\leq 2C^{\prime}(3\times 2^{-(k+4)}(m2^{k\gamma}-\mathfrak{z}))^{(2-\alpha)/(1-\alpha)}$. In this case, the last term in (37) is larger than
@@
Since $C_{k}\subset\{\mathscr{D}_{k}\neq\emptyset\}$, we have almost surely that, on $C_{k}$, $\{\min_{v\in\mathscr{U}_{k}^{m2^{k\gamma-1}}}Y^{m2^{k\gamma-1}}_{v}(t-)+\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2}-t);(t,i,\mathfrak{z})\in\mathscr{D}_{k}\}$ is a non-empty subset of the set of positions at time $2^{k+2}$ of particles in the system $Y^{m2^{k\gamma-1}}$ that are born in the time interval $2^{k},2^{k+1})$. In particular, we have almost surely that, on $C_{k}$,
$\min_{v\in\mathscr{U}_{k+1}^{m2^{k\gamma-1}}}Y^{m2^{k\gamma-1}}_{v}(2^{k+2})$ $\leq\min_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}\left\{\min_{v\in\mathscr{U}_{k}^{m2^{k\gamma-1}}}Y^{m2^{k\gamma-1}}_{v}(t-)+\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2}-t)\right\}$
$\leq\min_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}\left\{\min_{v\in\mathscr{U}_{k}^{m2^{k\gamma-1}}}Y^{m2^{k\gamma-1}}_{v}(2^{k+1})+\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2})\right\}$
$=\min_{v\in\mathscr{U}_{k}^{m2^{k\gamma-1}}}Y^{m2^{k\gamma-1}}_{v}(2^{k+1})+\min_{(t,i,\mathfrak{z})\in\mathscr{D}_{k}}\{\mathfrak{z}+\bm{\xi}_{t,i,\mathfrak{z}}(2^{k+2})\}$
$\leq\min_{v\in\mathscr{U}_{k}^{m2^{(k-1)\gamma-1}}}Y^{m2^{(k-1)\gamma-1}}_{v}(2^{k+1})+m2^{k\gamma}.$ (43)
@@
$\underline{Y}(s)\leq\underline{Y}(2^{k+1})\leq\min_{v\in\mathscr{U}_{k}^{m2^{(k-1)\gamma-1}}}Y^{m2^{(k-1)\gamma-1}}_{v}(2^{k+1})\leq\min_{v\in\mathscr{U}_{K_{0}}^{m2^{(K_{0}-1)\gamma-1}}}Y_{v}^{m2^{(K_{0}-1)\gamma-1}}(2^{K_{0}+1})+\sum_{j=K_{0}}^{k-1}m2^{j\gamma}.$
@@
$\sup_{s\in[2^{k},2^{k+1}]}\frac{\underline{Y}(s)}{s^{\gamma}}\leq\frac{m}{2^{\gamma}-1}+2^{-k\gamma}\left(\min_{v\in\mathscr{U}_{K_{0}}^{m2^{(K_{0}-1)\gamma-1}}}Y_{v}^{m2^{(K_{0}-1)\gamma-1}}(2^{K_{0}+1})-\frac{m2^{K_{0}\gamma}}{2^{\gamma}-1}\right).$